# Automatic Combination of Sample Selection Strategies for Few-Shot Learning

### 05 Feb 2024 v1

@@ Abstract
In few-shot learning, such as meta-learning, few-shot fine-tuning or in-context learning, the limited number of samples used to train a model have a significant impact on the overall success. Although a large number of sample selection strategies exist, their impact on the performance of few-shot learning is not extensively known, as most of them have been so far evaluated in typical supervised settings only. In this paper, we thoroughly investigate the impact of 20 sample selection strategies on the performance of 5 few-shot learning approaches over 8 image and 6 text datasets. In addition, we propose a new method for automatic combination of sample selection strategies (ACSESS) that leverages the strengths and complementary information of the individual strategies. The experimental results show that our method consistently outperforms the individual selection strategies, as well as the recently proposed method for selecting support examples for in-context learning. We also show a strong modality, dataset and approach dependence for the majority of strategies as well as their dependence on the number of shots – demonstrating that the sample selection strategies play a significant role for lower number of shots, but regresses to random selection at higher number of shots.
[…]
@@ 1 Introduction
Many domains are characterised by a labelled data scarcity due to data collection and annotation costs and privacy, making the use of typical deep learning unfeasible. To address these problems, few-shot learning aims to train or adapt machine learning models to new tasks using only few labelled samples per class (Song et al., 2023). There are two main categories of few-shot learning (more details regarding their difference are included in Appendix B): 1) gradient-based, which includes meta-learning and fine-tuning, where the knowledge from relevant datasets is transferred to new problem using gradient-based training, similar to supervised learning (Song et al., 2023; Chen et al., 2019; Vanschoren, 2018; Hospedales et al., 2021); and 2) few-shot learning via in-context learning, a new paradigm, where a pretrained large language model is conditioned on few examples to generate the prediction without any parameter updates (Dong et al., 2022; Liu et al., 2023).
As the number of utilised labelled samples is small, the role of individual samples is all the more important. Based on the choice of samples for in-context learning, the observed performance can vary between state-of-the-art and random predictions, indicating a strong sensitivity to the sample selection (Pecher et al., 2023; Köksal et al., 2022; Zhang et al., 2022). Even though a lot of research is dedicated to sample selection, such as active learning (Ren et al., 2021) or core-set selection (Guo et al., 2022), the most common approach in few-shot learning is a simple random selection. The exception is in-context learning, where a large number of sample selection strategies have been proposed recently, selecting samples according to their properties such as similarity, diversity, informativeness or quality (Li & Qiu, 2023; Zhang et al., 2022; Chang & Jia, 2023).
However, the impact of sample selection strategies on in-context learning or gradient-based few-shot learning remains largely under-explored. Therefore, the goal in this paper is to investigate the impact of the sample selection strategies on few-shot image and text classification tasks across a wide range of datasets.
@@
In addition, as the different strategies mostly focus on selecting samples based on a single property (e.g., how often the samples are forgotten (Toneva et al., 2018), how easy they are to learn (Swayamdipta et al., 2020; Zhang & Plank, 2021) or how informative they are (Ye et al., 2023b; Liu & Wang, 2023)), we ask whether the selection based on their combination would not lead to better results. To answer this question and to leverage the strengths of different sample selection strategies, we propose ACSESS, an effective method for Automatic Combination of SamplE Selection Strategies. The proposed method first identifies a subset of relevant strategies that can improve the overall success of few-shot learning. Afterwards, the identified strategies are combined together using weighting based on their expected contribution to find samples that can provide the most benefit. We summarise our contributions as follows111To support replicability and extension of our results, we openly publish the source code of our experiments at https://anonymous.4open.science/r/ACSESS:
@@
We evaluate the impact of 20 sample selection strategies on 5 few-shot learning approaches, including meta-learning, few-shot fine-tuning and in-context learning, across 8 image and 6 text datasets. Experimental results show the selection strategies often lead to non-negligible improvements in performance (up to percentage points for gradient few-shot learning or percentage points for in-context learning), while showing a strong approach and dataset dependence.
@@
We propose a method for automatic combination of sample selection strategies to leverage their strengths and identify samples with complementary properties required for the success of few-shot learning. Experimental results show the method consistently leads to performance improvement (up to percentage points for gradient few-shot learning or percentage points for in-context learning), outperforming all investigated strategies, including a recent in-context specific strategy LENS (Li & Qiu, 2023).
@@
We further analyse how the impact of sample selection changes when increasing the number of shots. We find following key insights: 1) sample selection has significantly higher impact on lower number of shots; 2) at higher number of shots (30-40 on average), the impact of sample selection is negligible as all strategies regress to random selection; and 3) after a certain point, the boost in performance from increasing the number of shots is negligible (50-shot for gradient few-shot learning; 20-shot for in-context learning with strong dependence on input text size and model context size).
@@ 2 Related Work
A lot of focus is dedicated to sample selection: 1) to reduce the computation costs without the loss of performance, by distilling dataset to lower number of samples (Yu et al., 2023); 2) selecting a core set of a dataset with similar properties and same performance as the full set (Guo et al., 2022); or 3) to reduce annotation costs using active learning (Ren et al., 2021). A specific research focus is dedicated to selecting good performing samples for in-context learning, as the overall performance was found to be sensitive to this choice, leading to large variability in results (Pecher et al., 2023; Köksal et al., 2022; Zhang et al., 2022). However, the adoption of sample selection in few-shot learning is limited to active learning strategies and in-context learning.
The sample selection strategies can be divided into two groups. The first, more prevalent, uses heuristics and unsupervised metrics to rate and select samples according to specific property that serves as estimate of their potential to increase the model’s performance. The most popular approach is selecting samples based on the similarity to the test sample for in-context learning. The methods differ in employed: 1) similarity measure, such as cosine/Euclidean distance (Liu et al., 2022; An et al., 2023; Zemlyanskiy et al., 2022; Gupta et al., 2022; Pasupat et al., 2021), BM-25 (Wang et al., 2022; Nashid et al., 2023) or their combination (Agrawal et al., 2023; Gao et al., 2021); and 2) representation, such as simple TF-IDF (Agrawal et al., 2023; Wang et al., 2022; Zemlyanskiy et al., 2022), word embeddings from basic or further fine-tuned transformers (Liu et al., 2022; Gao et al., 2021; Gupta et al., 2022; Pasupat et al., 2021), or representation generated by prompting a large language model (An et al., 2023; Nashid et al., 2023).
Besides similarity, the samples are often selected based on their informativeness or the uncertainty. The active learning strategies are most common (Köksal et al., 2022; Schröder et al., 2022; Margatina et al., 2021; Park et al., 2022; Mavromatis et al., 2023), but also other approaches are utilised, such as using the bias of the samples (Ma et al., 2023). Other methods define a notion of quality for each sample, either by prompting a large language model to rate the samples (Shin et al., 2021) or observing how the inclusion or removal of the sample affects the performance (Maronikolakis et al., 2023; Nguyen & Wong, 2023).
When considering representativeness, the samples are selected using core-set methods that are designed to find a subset of samples with properties that are representative of the full training set (Guo et al., 2022; Killamsetty et al., 2021; Mirzasoleiman et al., 2020; Paul et al., 2021; Coleman et al., 2020). Learnability metric either determines how easy it is to learn the specific sample (Swayamdipta et al., 2020; Zhang & Plank, 2021), or how often the sample is forgotten after being learned (Toneva et al., 2018).
@@
Finally, some approaches balance multiple properties at the same time. Often, the similarity of samples is balanced with their diversity (Wu et al., 2023; Qin et al., 2023) or the informativeness of the samples is balanced with their representativeness (Su et al., 2022; Levy et al., 2023; Ye et al., 2023b; Liu & Wang, 2023). In specific cases, the search is done in two step fashion, such as finding set of informative samples and then using diversity guided search to improve this set (Li & Qiu, 2023).
@@
The second set of approaches leverages training to select the samples. One possibility is to train a sample retriever, such as a separate scoring model (Rubin et al., 2022; Luo et al., 2023; Li et al., 2023; Wang et al., 2023; Aimen et al., 2023) or using determinantal point process (Yang et al., 2023; Ye et al., 2023a). Another set of approaches leverages reinforcement learning to select good samples (Zhang et al., 2022; Shum et al., 2023; Scarlatos & Lan, 2023). Finally, the Datamodels approach trains a linear model to predict the performance gain of a set of samples to select the subset that would lead to the highest possible performance increase (Ilyas et al., 2022; Chang & Jia, 2023; Jundi & Lapesa, 2022; Vilar et al., 2023).
In this work, we investigate the impact of the existing strategies on few-shot setting. In contrast to (Li & Qiu, 2023), which represents a work closest to ours, we explore multiple few-shot learning approaches not only in-context learning. Our proposed method is inspired by the Datamodels approach (Ilyas et al., 2022), but performs the selection at the level of strategies instead of samples.
3 Sample Selection Strategies
3.1 Single-Property Strategies
| Strategy | Description | |
Informativeness |
Similarity |
Most similar samples (popular for in-context learning). |
Diversity |
Most diverse/dissimilar samples. |
Entropy (Park et al., 2022) |
Active Learning; Highest entropy over class probabilities. |
Margin (Park et al., 2022) |
Active Learning; Lowest difference in probability between top 2 classes. |
Least Confidence (Park et al., 2022) |
Active Learning; Top class with lowest probability assigned. |
Loss (Park et al., 2022) |
Active Learning; Highest loss. |
Contrastive Active Learning (CAL) (Margatina et al., 2021) |
Core-Set; Predictive likelihood that diverges most from the neighbourhood as determined by KL divergence. |
GraNd (Paul et al., 2021) |
Core-Set; Highest contribution to decline of training loss. |
Core-Set; Submodularity function that measures diversity and informativeness. |
Representativeness |
Core-Set; Minimising distance of subset centre to full dataset. |
Core-Set; Minimising distance of every sample in subset to full dataset. |
CRAIG (Mirzasoleiman et al., 2020) |
Core-Set; Gradients representative of full dataset. |
Glister (Killamsetty et al., 2021) |
Core-Set; Bi-level optimisation. |
Learnability |
Forgetting (Toneva et al., 2018) |
Core-Set; How often the samples are forgotten. |
How easy it is to learn the samples. Due to no obvious consensus, all of easy, ambiguous, hard and combination of easy and ambiguous samples are considered. |
In this work, we thoroughly investigate the impact of strategies that consider 3 groups of properties: 1) informativeness, or how informative the samples are for the model; 2) representativeness, or how well the subset of samples represents the full dataset; and 3) learnability, or how easy it is to obtain or retain the information contained in the samples. We specifically focus on strategies that perform the selection by assigning a score to each sample that is used to sort the samples and choose the best. This includes active learning strategies that measure the uncertainty of the model, core-set selection strategies that find subset representative of the full-set and other strategies that use training dynamics. The high-level overview of these strategies is presented in Table 1 and their further details are included in Appendix C.
3.2 ACSESS: Automatic Combination of Strategies
@@
Even though the different selection strategies can identify samples with specific single property, the use of these identified samples may not have a significant impact on the performance when working in a few-shot setting. For example, an informative sample that is a hard to learn or is often forgotten, may not contribute as much, leading to low impact on the performance. Instead, we propose to select examples that are characterised by a balance of complementary properties, even when they do not represent the best sample of any given property (e.g., a sample that may be less informative, but is more easily learned or retained). To accomplish this, we propose ACSESS, an effective method for Automatic Combination of SamplE Selection Strategies, composed of two stages: 1) identification of a subset of well-performing and relevant strategies that can improve the overall success of few-shot learning; and then 2) combination of the identified strategies using a weighted scheme based on their expected contribution. The final set of samples is selected by sorting the samples according to their final score and selecting the ones with the highest score.
@@
In the first phase, the aim is to identify relevant strategies that, when combined, can select high quality samples based on their different properties. Even though the most optimal combination would be achieved by searching through all the combinations of strategies, such search is infeasible as it requires extensive computation costs. Instead, we use three separate methods for selecting strategies, inspired by feature selection in traditional machine learning and the Datamodels approach (Ilyas et al., 2022), specifically: 1) forward strategy selection; 2) backward strategy selection; and 3) datamodels strategy selection. The final subset of selection strategies is composed of the ones that are identified by each of the methods as most impactful, to minimise the effect of potential drawbacks of the individual method.
@@
Forward Selection iteratively adds to the subset the selection strategy that results in the highest performance increase (similarly to forward feature selection). Starting with an empty subset, the method alternates between two steps. First, for each strategy not included in the subset we determine the increase in performance the strategy would yield if added to the subset (evaluating the performance of samples selected by the combination of already selected subset and the potential new strategy and comparing to old subset). Afterwards, the strategy that provides the highest positive increase in performance is added to the subset and the first step is repeated. If no added strategy leads to increase of the overall performance, the process ends, resulting in the subset of relevant selection strategies.
@@
Backward Selection iteratively removes strategies from the subset until the performance no longer increases (similarly to backward feature selection). The method works the same way as the forward selection, but starting with all selection strategies and removing them from the subset based on the performance increase such removal leads to.
@@
Datamodels Selection draws inspiration from the Datamodels (Ilyas et al., 2022; Chang & Jia, 2023), but works on the level of the strategies instead of samples. First, a set of 150 random combinations of strategies is created and evaluated. The random search is constrained to include each strategy at least 5 times to ensure sufficient coverage of each strategy. Afterwards, these combinations are used to create a simple regression dataset, with the difference in performance to the baseline selection (the classic few-shot learning selection) as label, and the multi-label presence vector of strategies as features. The dataset is used to train a regularised linear regression, LASSO (Tibshirani, 1996), which zeros out the weights for as many strategies as possible. The selection strategies with positive weights in the model represent the subset of relevant strategies.
@@
After the relevant selection strategies are identified, they are combined using a weighting scheme. Each sample is assigned an importance score by each of the identified strategies. As the different strategies may assign scores of different magnitudes, all scores are normalised to zero-one interval. The final sample score is obtained as a weighted average of its assigned scores and the samples with highest score is selected.
In this paper, we evaluate three separate weighting schemes. First, uniform combination, assigns the same weight to each of the identified strategies. This represents the simplified version of the ACSESS method, potentially resulting in lower performance impact, but also with lower computation costs as no further optimisation is required. This weighting is also used for forward and backward selection. Second, weighted combination, assigns a different weight to each strategy, prioritising specific sample properties. The final trained weights of the Datamodels linear model are used as the dataset and model specific weights for the different strategies. As the weighting scheme is dataset and model specific, it should lead to the highest performance impact, at the cost of requiring more extensive computation. Finally, the weighted combination with random selection introduces a random element to the weighted combination, as random selection was identified as a strong baseline that can improve active learning performance (Park et al., 2022). Each sample is assigned additional uniformly randomised score which is included in the combination. The weight of this score is determined by the Datamodels by including random selection as additional strategy, due to its impact being dataset and model dependent.

### 17 Apr 2026 v2

@@ Abstract
In few-shot learning, the selection of samples has a significant impact on the performance of the model. While effective sample selection strategies are well-established in supervised settings, research on large language models largely overlooks them, favouring strategies specifically tailored to individual in-context learning settings. In this paper, we propose a new method for Automatic Combination of SamplE Selection Strategies (ACSESS) to leverage the strengths and complementarity of various well-established selection objectives. We investigate and compare the impact of 23 sample selection strategies on the performance of 5 in-context learning models and 3 few-shot learning approaches (meta-learning, few-shot fine-tuning) over 6 text and 8 image datasets. The experimental results show that the combination of strategies through the ACSESS method consistently outperforms all individual selection strategies and performs on par or exceeds the in-context learning specific baselines. Lastly, we demonstrate that sample selection remains effective even on smaller datasets, yielding the greatest benefits when only a few shots are selected, while its advantage diminishes as the number of shots increases.
Automatic Combination of Sample Selection Strategies for Few-Shot Learning
Branislav Pecher♠†, Ivan Srba†, Maria Bielikova†, Joaquin Vanschoren‡ ♠ Faculty of Information Technology, Brno University of Technology, Brno, Czechia † Kempelen Institute of Intelligent Technologies, Bratislava, Slovakia ‡ Eindhoven University of Technology, Eindhoven, Netherlands {branislav.pecher, ivan.srba, maria.bielikova}@kinit.sk, j.vanschoren@tue.nl
@@ 1 Introduction
Many domains are characterised by a labelled data scarcity due to data collection/annotation costs or privacy considerations, making the training of typical deep learning models unfeasible. Few-shot learning addresses the challenge of adapting models to new tasks when only a handful of labelled samples are available (Song et al., 2023). Two main approaches exist: 1) in-context learning, where a pretrained large language model is conditioned on few examples without any parameter updates (Dong et al., 2022; Liu et al., 2023); and 2) gradient-based few-shot learning, which includes meta-learning and fine-tuning, where the knowledge from relevant datasets is transferred to a new problem using gradient-based training (Song et al., 2023; Chen et al., 2019; Vanschoren, 2018; Hospedales et al., 2021). More details regarding their differences are included in Appendix J.
The choice of which sample to use is critical in both approaches, as performance can vary drastically depending on sample quality. Existing studies in different settings, such as in LLM alignment (Zhou et al., 2024), have shown that curating a smaller subset of high-quality samples can lead to better performance than training on the full set of available samples. This holds for few-shot learning as well, especially for noisy or unbalanced datasets or for in-context learning, which is known to be sensitive to sample choice (Pecher et al., 2024; Agarwal et al., 2021; Köksal et al., 2022).
Many works have already proposed diverse sample selection strategies to identify an optimal set of samples. Such a sample selection strategy follows its specific objective by considering one or few properties of available samples, e.g., to maximise similarity, diversity, informativeness or quality of selected samples (Li and Qiu, 2023; Zhang et al., 2022; Chang and Jia, 2023). With the popularisation of large language models (LLMs), many new selection strategies are proposed for in-context learning. However, they are often tailored only for specific settings, limiting their applicability. At the same time, the well-established and long-standing selection strategies that were shown to perform well in supervised settings, such as active learning, are mostly ignored as part of LLM sample selection (Pecher et al., 2024; Albalak et al., 2024).
@@
In this work, we demonstrate that well-established strategies can match or exceed the performance of in-context learning specific strategies. To achieve this, we automatically and optimally combine the sample selection objectives used by well-established sample selection strategies to identify samples with complementary properties. In order to leverage the strengths of different sample selection strategies, we propose ACSESS, an effective method for Automatic Combination of SamplE Selection Strategies. First, a subset of relevant strategies that can improve the overall success of few-shot learning is identified. Afterwards, the selection objectives from the identified strategies are combined together, using weighting based on their expected contribution, in order to identify the most informative and high-quality samples that can provide the most benefit. Our contributions and findings are as follows111To support replicability and extension of our results, we openly publish the source code of our experiments at https://github.com/kinit-sk/ACSESS:
@@
We provide the first large-scale study of 23 sample selection strategies across 5 in-context learning models and 3 few-shot learning approaches, including meta-learning and few-shot fine-tuning, across 6 text and 8 image datasets. We show that sample selection can yield consistent gains (up to and percentage points for in-context learning and gradient few-shot learning, respectively), though the effects depend heavily on dataset and approach.
@@
We propose ACSESS, a method that automatically combines and weights sample selection objectives to leverage their strengths and identify samples with complementary properties (i.e., informativeness, representativeness and learnability). Our results show that the combination of well-established strategies identified by the ACSESS method consistently leads to performance on par or better than the existing in-context learning specific baselines.
@@
Through ablation studies, we find following key insights: 1) sample selection has higher impact when the number of shots is low or when using noisy datasets; 2) at higher number of shots (30-40 on average), the impact of sample selection is negligible as all strategies regress to random selection; 3) after a certain point, the boost in performance from using more shots becomes negligible (30 shots for in-context learning; 50-shot for gradient few-shot learning); 4) sample selection is beneficial even for small dataset sizes (achieving similar performance when selecting from only 25% for in-context learning or 10% of the dataset for gradient few-shot learning).
@@ 2 Related Work
A large body of work is dedicated to selecting high-quality samples for in-context learning, where the overall performance was found to be sensitive to this choice, leading to large variability in results (Pecher et al., 2024; Köksal et al., 2022; Zhang et al., 2022). In the supervised setting, the selection is mostly focused on distilling datasets to a lower number of samples (Yu et al., 2023), selecting a core set representative of the full set (Guo et al., 2022), or reducing annotation costs using active learning (Ren et al., 2021). However, the impact of these strategies remains underexplored for other few-shot learning approaches.
Within the majority of the sample selection strategies, the sample selection objective takes into consideration a single property of available samples. To this end, various heuristics and unsupervised metrics serve as an estimate of their potential to increase the model’s performance. For in-context learning, the most popular approach is selecting samples based on the similarity to the test sample, differing only in employed similarity measure and representation (Liu et al., 2022; An et al., 2023; Zemlyanskiy et al., 2022; Gupta et al., 2022; Pasupat et al., 2021; Wang et al., 2022; Nashid et al., 2023; Agrawal et al., 2023; Gao et al., 2021). Besides similarity, the samples are often selected based on their informativeness or the uncertainty, often using active learning strategies (Köksal et al., 2022; Schröder et al., 2022; Margatina et al., 2021; Park et al., 2022; Mavromatis et al., 2023), bias of the samples (Ma et al., 2023) or adversarial training (Agarwal et al., 2021). Other methods define a notion of quality for each sample, either by prompting a LLM to rate the samples (Shin et al., 2021) or observing how the inclusion or removal of the sample affects the performance (Maronikolakis et al., 2023; Nguyen and Wong, 2023). When considering representativeness, the samples are selected based on how well they are representative of the full dataset (Guo et al., 2022; Killamsetty et al., 2021; Mirzasoleiman et al., 2020; Paul et al., 2021; Coleman et al., 2020). Learnability metric either determines how easy it is to learn the specific sample (Swayamdipta et al., 2020; Zhang and Plank, 2021), or how often the sample is forgotten after being learned (Toneva et al., 2018).
@@
Finally, some approaches balance multiple sample properties at the same time. Often, the similarity of samples is combined with their diversity (Wu et al., 2023; Qin et al., 2023) or the informativeness of the samples is balanced with their representativeness (Su et al., 2022; Levy et al., 2023; Ye et al., 2023; Liu and Wang, 2023). In specific cases, a two step sample search is done, such as finding set of informative samples and then using diversity guided search to improve this set (Li and Qiu, 2023), or using different uncertainty measures to select the top samples (Diao et al., 2024).
@@
Newer strategies leverage optimisation to select a single representative set of samples. One possibility is to train a sample retriever, such as a separate scoring model (Rubin et al., 2022; Luo et al., 2023; Li et al., 2023; Wang et al., 2023; Aimen et al., 2023) or define a surrogate scorer for modelling the sample’s goodness and use multi-armed bandit to select the best exemplars (Purohit et al., 2024, 2025). Another option is to leverage reinforcement learning (Zhang et al., 2022; Shum et al., 2023; Scarlatos and Lan, 2023). Finally, the Datamodels approach trains a linear model to predict the performance gain of a set of samples to select the subset that would lead to the highest possible performance increase (Ilyas et al., 2022; Chang and Jia, 2023; Jundi and Lapesa, 2022; Vilar et al., 2023).
In this work, we investigate whether the combination of objectives from well-established single-property strategies can rival the selection from the strategies tailored specifically to in-context learning. Similar to the close works of Li and Qiu (2023); Purohit et al. (2024, 2025), we select a single set of samples, but do not focus solely on in-context learning. In essence, we complement work by Agarwal et al. (2021) by exploring and effectively combining sample properties important for few-shot learning performance. Finally, our proposed method is inspired by the Datamodels approach (Ilyas et al., 2022), but performs the selection at the level of strategies instead of samples.
3 ACSESS: Automatic Combination of Selection Strategies
Despite the large number of existing sample selection strategies, most remain constrained to a single objective, typically considering only one sample property. This narrow focus often leads to suboptimal selection, whereas greater performance gains can be achieved by appropriately combining multiple objectives that capture complementary sample properties. For example, the most informative sample that is hard to learn or is often forgotten may not contribute as much for few-shot methods that utilise gradient-based training. On the other hand, for in-context learning, the samples on the decision boundary, which are often hard to learn, may provide more benefit.
@@ 3 ACSESS: Automatic Combination of Selection Strategies
Instead of focusing on a single property, we propose to select examples that are characterised by a balance of complementary properties, even when they may not represent the best sample of any given individual property. To accomplish this, we propose ACSESS, a method for Automatic Combination of SamplE Selection Strategies taking their complementarity into account. The method is composed of three stages: 1) definition of single-property strategies that will be considered for the combination; 2) identification of a subset of well-performing and relevant strategies that can improve the performance of few-shot learning; and then 3) combining the identified strategies to identify the final set of the most beneficial samples.
To combine selection objectives, the ACSESS method assigns a selection score for each sample, representing its expected contribution towards a higher performance. It is calculated as a weighted sum of scores provided by individual objectives, formally , where represents the set of considered strategies, is the weight assigned to the strategy , is the input sample to be scored, and is the score assigned to the sample by strategy normalised to a interval. After assigning a score to each sample , samples with the highest score are selected. The steps of the method are summarised in Algorithm 1, with further details, including discussion of the computation complexity and its optimisation, in Appendix A.
Single-Property Strategies
We focus on well-established strategies that select based on following properties: 1) informativeness, or how informative the samples are for the model; 2) representativeness, or how well the subset of samples represents the full dataset; and 3) learnability, or how easy it is to obtain or retain the information contained in the samples. In this work, we specifically focus on strategies that work by assigning a score to each sample, sorting based on this score and selecting the samples with the best score. For informativeness, this includes selecting samples based in their Similarity and Diversity, active learning strategies, such as Entropy, Margin, Least Confidence and Loss (Park et al., 2022), and core-set selection including Contrastive Active Learning (CAL) (Margatina et al., 2021), DeepFool (Ducoffe and Precioso, 2018), GraNd (Paul et al., 2021) and Graph-Cut (Iyer and Bilmes, 2013; Iyer et al., 2021). For representativeness, we consider Herding (Welling, 2009; Chen et al., 2010), KCenter (Sener and Savarese, 2018; Agarwal et al., 2020), CRAIG (Mirzasoleiman et al., 2020) and Glister (Killamsetty et al., 2021). Finally, for learnability, we consider Forgetting (Toneva et al., 2018) and Cartography (Swayamdipta et al., 2020; Zhang and Plank, 2021), selecting all of each, ambiguous, hard and a combination of easy and ambiguous samples (due to no obvious consensus on which are the best). More details regarding the strategies and their objectives are included in Appendix B.
@@
In the second phase, the goal is to identify relevant strategies that, when combined, can select high-quality samples based on their complementary properties. Even though the optimal combination could be achieved by searching through all the combinations of strategies, such a search is infeasible as it requires extensive computation. Instead, we draw inspiration from methods commonly used for feature selection in traditional machine learning – namely the forward and backward selection – and for the first time apply them to the problem of sample selection. In addition, we adapt Datamodels, introduced by Ilyas et al. (2022) for individual sample selection, to work on the level of strategies instead. Such an approach allows us to explore a more diverse set of possible strategy combinations as compared to forward/backward selection. All three of the methods are run independently of each other, each producing a separate set of strategies, i.e., , and respectively for forward, backward and datamodels selection. The final set of strategies used by our method is constructed as an intersection between these sets, i.e., . This allows the method to identify the most impactful strategies (i.e., strategies identified by all of the methods), minimise the potential drawbacks of the individual methods (i.e., exploring also diverse combinations), and guarantee efficiency of the method, by identifying the lowest number of strategies possible.
@@
Forward Selection iteratively adds strategies that result in the highest performance increase (similarly to forward feature selection). Starting with an empty subset, the method alternates between two steps. First, for each strategy not included in the subset of selected strategies, we determine the increase in performance the strategy would yield if added to the subset. We determine the increase by training the model with samples with the highest selection score determined by so far selected strategies (with weights uniformly distributed for all strategies, ). Second, the strategy that provides the highest positive increase in performance is added to the subset and the first step is repeated. If the addition of any strategy does not lead to a performance increase, the process ends, resulting in the subset of relevant selection strategies () (see Algorithm 2 in Appendix A for more details). In essence, the forward selection iteratively adds the best-performing strategies to the set while the performance still increases.
@@
Backward Selection iteratively removes strategies until the performance no longer increases (similarly to backward feature selection). The method works the same way as the forward selection, but starts with all selection strategies and removes them from the subset () based on the performance increase such removal leads to (see Algorithm 3 in Appendix A for more details). In essence, the backward selection iteratively removes strategies that lead to lower performance until the performance no longer increases.
@@
Datamodels Selection draws inspiration from Datamodels (Ilyas et al., 2022; Chang and Jia, 2023). First, a set of (in our work ) random combinations of strategies is created and evaluated (using uniform combination). The random search is constrained to include each strategy at least 5 times to ensure sufficient coverage of each strategy. Afterwards, these combinations and their performance are used to create a simple regression dataset. The difference in performance to the baseline selection (the classic few-shot learning selection) is used as the target. As features, we use a multi-label presence vector of strategies, i.e., vector containing only zeros and ones, where 1 is assigned to elements corresponding to strategy that was used in the combination, and 0 for all other ones. The dataset is then used to train a regularised linear regression, LASSO (Tibshirani, 1996), which zeros out the weights for as many strategies as possible. The selection strategies with positive weights in the model represent the subset of relevant strategies () (see Algorithm 4 in Appendix A for more details).
@@
After identifying the relevant selection strategies (), a weighted combination of objectives is used to perform the selection. The final set consists of samples with the highest selection score.
In this paper, we evaluate three separate weighting schemes. First, uniform combination, assigns the same weight to each of the identified strategies (). This represents the simplified version of the ACSESS method that potentially leads to lower performance impact, but also lower computation costs and better transferability, as no further optimisation is required. This weighting scheme is also used for the previous step for the forward, backward and datamodels selection. Second, weighted combination, assigns a different weight to each strategy objective, prioritising specific sample properties beneficial for the specific few-shot approach. As the strategy weights for this weighted combination, we use the final trained weights of the Datamodels linear model. As this weighting scheme is dataset and model-specific, it often leads to the highest performance impact, at the cost of requiring more extensive computation. For this reason, in this paper, we evaluate the strategy on a more aggregated level in terms of datasets – see Appendix A for details. Finally, the weighted combination with random selection extends the weighted combination by introducing a random element to it, as the random selection was identified as a strong baseline (Park et al., 2022; Cegin et al., 2025). For this weighting scheme, each sample is assigned an additional uniformly randomised score. To determine the weight of this score, the random selection is included as an additional strategy to the Datamodels method.